OpenStudy (anonymous):

The given measurements may or may not determine a triangle. B = 137°, c = 6, b = 11

OpenStudy (anonymous):

The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used.

OpenStudy (anonymous):

@Loser66 @hartnn @amistre64 Anybody? :(

OpenStudy (campbell_st):

well looking at the information you know b is the longest side since its opposite the largest angle B. well the law of sines says \[\frac{\sin(A)}{a} = \frac{\sin(B)}{b} = \frac{\sin(C)}{c}\] you have B and b and c so \[\frac{\sin(137)}{11} = \frac{\sin(C)}{6}\] so angle C = 21.84 degrees. So by angle some of a triangle angle A = 21.16 find side a by the law of sines... and it should be smaller than side c \[\frac{11}{\sin(137)} = \frac{a}{\sin(21.16)}\] a = 5.82 now check that a + c > b 6+ 5.8 > 11 true so you have a triangle.

OpenStudy (anonymous):

Thanks campbell! So much help in just one reply